# CSI-BERT2: A BERT-inspired Framework for Efficient CSI Prediction and Classification in Wireless Communication and Sensing

### v3

@@ Abstract
Channel state information (CSI) is a fundamental component in both wireless communication and sensing systems, enabling critical functions such as radio resource optimization and environmental perception. In wireless sensing, data scarcity and packet loss hinder efficient model training, while in wireless communication, high-dimensional CSI matrices and short coherent times caused by high mobility present challenges in CSI estimation. To address these issues, we propose a unified framework named CSI-BERT2 for CSI prediction and classification tasks, built on CSI-BERT, which adapts BERT to capture the complex relationships among CSI sequences through a bidirectional self-attention mechanism. We introduce a two-stage training method that first uses a mask language model (MLM) to enable the model to learn general feature extraction from scarce datasets in an unsupervised manner, followed by fine-tuning for specific downstream tasks. Specifically, we extend MLM into a mask prediction model (MPM), which efficiently addresses the CSI prediction task. To further enhance the representation capacity of CSI data, we modify the structure of the original CSI-BERT. We introduce an adaptive re-weighting layer (ARL) to enhance subcarrier representation and a multi-layer perceptron (MLP)-based temporal embedding module to mitigate temporal information loss problem inherent in the original Transformer. Extensive experiments on both real-world collected and simulated datasets demonstrate that CSI-BERT2 achieves state-of-the-art performance across all tasks. Our results further show that CSI-BERT2 generalizes effectively across varying sampling rates and robustly handles discontinuous CSI sequences caused by packet loss—challenges that conventional methods fail to address. The dataset and code are publicly available at https://github.com/RS2002/CSI-BERT2.
@@ I Introduction
Channel state information (CSI) plays a critical role in both wireless sensing and wireless communication systems by capturing the propagation characteristics of the wireless channel. This information provides valuable insights into the radio environment, including the channel conditions, signal quality, and the motion and location of objects. In wireless communication systems, CSI is essential for optimizing key processes in wireless communication systems, such as channel compensation, adaptive modulation and coding, user selection and scheduling. Effective utilization of CSI enables improved system performance, increased spectrum efficiency, and more reliable communication, ultimately enhancing the overall user experience [1, 2, 3]. In wireless sensing applications, CSI can enhance the ability to detect and track objects, leading to improved environmental awareness [4, 5, 6]. Owing to its advantages in privacy protection, low cost, and penetration ability, wireless sensing has been widely adopted in diverse applications, such as fall detection [7], localization [8], gesture recognition [9], and person identification [10].
@@
Despite the importance of CSI in wireless systems, acquiring, processing, and utilizing CSI data poses significant challenges. In wireless communication, estimating CSI is time-consuming, particularly in high-dimensional and high-mobility scenarios. Frequent updates in rapidly changing environments increase computational complexity. Consequently, delays in obtaining accurate CSI can degrade spectrum efficiency by leading to suboptimal resource allocation and increased interference. In wireless sensing, packet loss due to factors such as channel noise and device errors may severely impact model performance. As noted in [11], even in relatively simple indoor environments, the average packet loss rate can reach 14.5%, with peaks up to 70% within a single second. Such discontinuities in CSI sequences pose serious challenges for the design and deployment of wireless sensing models. Additionally, for data-driven methods in both scenarios, a major challenge is the limited availability of datasets. Collecting large datasets independently is costly, while public datasets are often hetergeneous and scarce, making them difficult to use in combination. Consequently, how to enable models to learn effectively from limited datasets has become an important research topic.
@@
Recent studies have noted and addressed these challenges. In wireless communication, the high computational complexity of CSI estimation has driven the adoption of deep learning techniques to predict future CSI more efficiently. However, most existing approaches rely on auto-regressive (AR) manners [12, 13, 14, 15], which have been criticized for error accumulation [16]. Regarding the packet loss problem in wireless sensing, most previous methods employ interpolation techniques such as linear interpolation and Kriging to estimate CSI within each dimension, often neglecting the inter-dependencies among subcarriers. Prior research [11] has indicated that these interpolation methods provide only limited benefits for sensing models. To address the issue of data scarcity, the pre-training and fine-tuning paradigm has been extensively explored in wireless systems and has shown promising performance [17]. By pre-training on either downstream datasets or unlabeled datasets, the model can achieve a more generalized capacity for data representation, thereby enhancing its generalization and robustness in downstream tasks. [^1] To avoid confusion, we clarify the concept of “pre-training” in this paper. We use this term in its original and broadest sense: any initial training phase that improves model performance on downstream tasks, regardless of the data source (labeled/unlabeled, in-domain/out-of-domain) or methodology (self-supervised, supervised, etc.). This intentionally differs from narrower interpretations that restrict “pre-training” exclusively to large-scale unsupervised learning on massive datasets, as commonly employed in today’s large models. Finally, given the wide diversity of tasks in wireless communication and sensing, designing specialized models for each task can be labor-intensive and inefficient. A unified and generic framework is therefore highly desirable [18], and many works [19, 20, 21] have begun exploring the development of foundational models for wireless systems.
@@
To address the aforementioned shortcomings, we propose a method inspired by bidirectional encoder representations from transformers (BERT) [22], named CSI-BERT2, which is an improved version of our previous work, CSI-BERT [11]. In CSI-BERT, the masked language model (MLM) approach was first introduced for the pre-training of the CSI recovery model in Wi-Fi sensing, enabling it to learn effectively from limited dataset and improving performance in downstream tasks. Building on this, CSI-BERT2 introduces the mask prediction model (MPM), where we mask the final portion of the CSI sequence and train the model to recover it. By predicting the masked CSI in a single step, our method efficiently avoids the error accumulation problem common in AR models. Additionally, we propose a novel multi-layer perceptron (MLP)-based time embedding method that effectively addresses the temporal information loss problem [23] inherent in the original Transformer. As illustrated in [23], the permutation-invariant nature of the attention mechanism leads to temporal information loss, even with positional embeddings. In contrast, our proposed time embedding method explicitly exploits the temporal order, which is critical for CSI sequences, thereby mitigating this issue. This design enables CSI-BERT2 to generalize efficiently in environments with varying sampling rates, and perform robustly even in a zero-shot setting. Finally, considering the varying information content convened across different subcarriers, we introduce an adaptive re-weighting layer (ARL) [24], which helps to adaptively capture the importance of each subcarrier. Through these architectural enhancements, we successfully address the low performance of the original CSI-BERT in CSI classification tasks. In conclusion, the main contributions of this paper are summarized as follows:
@@ II-A CSI Basics
Following [25], in Wi-Fi communication, CSI characterizes the way Wi-Fi signals propagate from the transmitter (TX) to the receiver (RX). Modern Wi-Fi systems typically employ orthogonal frequency division multiplexing (OFDM) as the modulation technique. OFDM divides the wideband channel into multiple narrowband subcarriers in the frequency domain, each experiencing flat fading for easier processing. In the time domain, data is transmitted as a series of OFDM symbols with a cyclic prefix to prevent interference from multipath delays. The amplitude and phase of CSI are influenced by multipath effects, which include amplitude attenuation and phase shift. Each element of the CSI matrix $H$ is given by:
@@ II-B CSI Prediction
Obtaining accurate CSI matrix in wireless communication remains a challenging task [26]. The large number of antennas results in a high-dimensional CSI matrix, which in turn makes the channel estimation algorithms difficult to solve. Additionally, when objects in the environment move at high speeds, the channel coherence time shortens considerably, imposing stricter requirements on the channel estimation speed. As illustrated in Fig. 2b, CSI plays a critical role in optimizing wireless communication system performance, making fast and reliable CSI estimation essential. Furthermore, predicting future CSI can provide valuable opportunities for system-level optimizations, such as effective channel compensation.
Recently, some works have explored the use of additional information for CSI prediction. For example, in [27], the authors used multimodal sensory data, including received pilots, user position, and historical downlink CSI to predict the current CSI. In [28], a deep learning-based framework using a three dimensional (3D) CNN architecture effectively predicts downlink CSI by learning temporal, spatial, and frequency correlations from past channel sequences. The authors in [29] have proposed an machine learning (ML)-based architecture for estimating channels by using CNN-AR, which achieves gains in prediction quality.
@@ II-C CSI Classification
In wireless sensing, many tasks can be formulated as CSI classification problems, including gesture recognition[30], fall detection [7], and activity recognition [31]. Reviewing CSI basics, human movement in the environment can cause changes in the multipath channel, which in turn changes the extracted CSI matrix. Therefore, by analyzing the pattern of changes in CSI, it is possible to sense human activities.
However, in real-world sensing systems, accurately characterizing human actions are challenging. Complex actions, such as falling, kicking, and squatting, are particularly difficult to model mathematically. Additionally, noise in Wi-Fi extraction devices, including carrier frequency offset and sampling time offset, is hard to completely eliminate. These challenges motivate the use of deep learning techniques for Wi-Fi sensing. For example, in [32], the authors have estimated human poses from videos to supervise the training of Wi-Fi CSI. The broad learning system proposed in [33] aims to achieve fast and accurate Wi-Fi fingerprint localization. Furthermore, [34] have proposed Wi-Mesh, which utilizes Wi-Fi signals and 2D angle-of-arrival (AoA) estimation to create 3D human meshes using deep learning, showing robustness across various environments and conditions.
@@
To tackle this challenge, our previous work have introduced CSI-BERT [11], a self-supervised learning method designed to capture the internal structure of CSI data and recover lost packets. The recovered CSI was shown to substantially enhance the performance of downstream sensing tasks. In this work, we further optimize the structure of CSI-BERT to improve its ability to capture temporal information, enabling it to successfully perform CSI classification tasks using the discontinuous CSI for training.
@@ III-A Overview
In this section, we will introduce the data preprocessing methods, neural network architecture, pre-training process, and the fine-tuning and inference methods of CSI-BERT2 in turn. As a continuation of CSI-BERT, we first provide a brief introduction to it. (This section will primarily focus on the modified and enhanced components, while more details about CSI-BERT can be found in our previous paper [11].)
Prior to CSI-BERT, several works had already applied BERT [22] to wireless sensing tasks such as people localization [36, 37] and radio map construction [38], based on the intuition that the bidirectional encoder of the Transformer [39] can effectively capture the relationships among packets by the self-attention mechanism. However, these studies typically fed discrete signals into the original BERT embedding layer, which leads to significant information loss. In contrast, CSI-BERT first proposed a novel embedding layer tailored to the structure of CSI data, where a multi-layer perceptron (MLP) is used to encode continuous CSI data directly. Additionally, a positional-embedding style time-embedding layer was introduced to capture the temporal information among packets. Inspired by the similarity between packet recovery and masked language modeling (MLM), CSI-BERT successfully achieves high-quality packet recovery through unsupervised learning without labeled data. The recovered CSI data enhances the performance of other sensing models in classification tasks, outperforming traditional interpolation methods that fail to consider the internal relationships among subcarriers.
However, the original CSI-BERT does not fully leverage the advantages of MLM, resulting in subpar performance in downstream tasks. For example, the classification accuracy of CSI-BERT after fine-tuning does not surpass that of the vanilla ResNet18 [40]. This limitation can be attributed to the inadequacies of the embedding layer. To address this issue, CSI-BERT2 introduces an ARL-based spatial encoding module to capture the relationships and importance among subcarriers. Additionally, a MLP-based temporal encoding module is proposed to capture both the order and distance relationships among packets, effectively solving the temporal information loss problem inherent in conventional positional embedding modules, which capture distance relationships without considering order information. Moreover, to fully utilize the learned knowledge during pre-training, we propose a novel fine-tuning method named mask prediction model (MPM) for the CSI prediction task, wherein the last portion of the CSI sequence is randomly masked, and the model is trained to predict these masked values. Compared to conventional AR-based methods, our approach can complete the prediction in one step, thereby eliminating the error accumulation problem.
@@
TABLE I: Comparison between CSI-BERT [11] and CSI-BERT2: Here PE represents the original positional embedding in Transformer, and SFT represents supervised fine-tuning. We utilize underline to highlight the differences between CSI-BERT2 and CSI-BERT.
CSI-BERT [11] CSI-BERT2 Model Structure Spatial Embedding MLP MLP + ARL Temporal Embedding PE PE + MLP Positional Embedding PE PE Backbone BERT BERT Discriminator DANN-based GAN-based Training Method Pre-training MLM MLM Classification Fine-tuning SFT SFT + Random Mask Prediction Fine-tuning $\times$ MPM Application Scenarios CSI Recovery $\checkmark$ $\checkmark$ CSI Classification Single Sampling Rate Various Sampling Rates CSI Prediction $\times$ $\checkmark$

### v4

@@ Abstract
Channel state information (CSI) is a fundamental component in both wireless communication and sensing systems, enabling critical functions such as radio resource optimization and environmental perception. In wireless sensing, data scarcity and packet loss hinder efficient model training, while in wireless communication, high-dimensional CSI matrices and short coherent times caused by high mobility present challenges in CSI estimation. To address these issues, we propose a unified framework named CSI-BERT2 for CSI prediction and classification tasks, built on our previous work CSI-BERT, which adapts BERT to capture the complex relationships among CSI sequences through a bidirectional self-attention mechanism. We introduce a two-stage training method that first uses a mask language model (MLM) to enable the model to learn general feature extraction from scarce datasets in an unsupervised manner, followed by fine-tuning for specific downstream tasks. Specifically, we extend MLM into a mask prediction model (MPM), which efficiently addresses the CSI prediction task. To further enhance the representation capacity of CSI data, we modify the structure of the original CSI-BERT. We introduce an adaptive re-weighting layer (ARL) to enhance subcarrier representation and a multi-layer perceptron (MLP)-based temporal embedding module to mitigate temporal information loss problem inherent in the original Transformer. Extensive experiments on both real-world collected and simulated datasets demonstrate that CSI-BERT2 achieves state-of-the-art performance across all tasks. Our results further show that CSI-BERT2 generalizes effectively across varying sampling rates and robustly handles discontinuous CSI sequences caused by packet loss—challenges that conventional methods fail to address. The dataset and code are publicly available at https://github.com/RS2002/CSI-BERT2.
@@ I Introduction
Channel state information (CSI) plays a critical role in both wireless sensing and wireless communication systems by capturing the propagation characteristics of the wireless channel. This information provides valuable insights into the radio environment, including the channel conditions, signal quality, and the motion and location of objects. In wireless communication systems, CSI is essential for optimizing key processes in wireless communication systems, such as channel compensation, adaptive modulation and coding, user selection and scheduling. Effective utilization of CSI enables improved system performance, increased spectrum efficiency, and more reliable communication, ultimately enhancing the overall user experience [zhu2023pushing, li2023integrated, wen2023task]. In wireless sensing applications, CSI can enhance the ability to detect and track objects, leading to improved environmental awareness [xu2023edge, yao2024wireless, yang2025privacy]. Owing to its advantages in privacy protection, low cost, and penetration ability, wireless sensing has been widely adopted in diverse applications, such as fall detection [cai2023falldewideo], localization [lofi], gesture recognition [zhao2025does], and person identification [CSi-Net].
@@
Despite the importance of CSI in wireless systems, acquiring, processing, and utilizing CSI data poses significant challenges. In wireless communication, estimating CSI is time-consuming, particularly in high-dimensional and high-mobility scenarios. Frequent updates in rapidly changing environments increase computational complexity. Consequently, delays in obtaining accurate CSI can degrade spectrum efficiency by leading to suboptimal resource allocation and increased interference. In wireless sensing, packet loss due to factors such as channel noise and device errors may severely impact model performance. As noted in [CSI-BERT], even in relatively simple indoor environments, the average packet loss rate can reach 14.5%, with peaks up to 70% within a single second. Such discontinuities in CSI sequences pose serious challenges for the design and deployment of wireless sensing models. Additionally, for data-driven methods in both scenarios, a major challenge is the limited availability of datasets. Collecting large datasets independently is costly, while public datasets are often hetergeneous and scarce, making them difficult to use in combination. Consequently, how to enable models to learn effectively from limited datasets has become an important research topic.
@@
Recent studies have noted and addressed these challenges. In wireless communication, the high computational complexity of CSI estimation has driven the adoption of deep learning techniques to predict future CSI more efficiently. However, most existing approaches rely on auto-regressive (AR) manners [LSTM_CSI1, LSTM_CSI2, LSTM_CSI3, LSTM_CSI4], which have been criticized for error accumulation [parthipan2024defining]. Regarding the packet loss problem in wireless sensing, most previous methods employ interpolation techniques such as linear interpolation and Kriging to estimate CSI within each dimension, often neglecting the inter-dependencies among subcarriers. Prior research [CSI-BERT] has indicated that these interpolation methods provide only limited benefits for sensing models. To address the issue of data scarcity, the pre-training and fine-tuning paradigm has been extensively explored in wireless systems and has shown promising performance [lyu2024rethinking]. By pre-training on either downstream datasets or unlabeled datasets, the model can achieve a more generalized capacity for data representation, thereby enhancing its generalization and robustness in downstream tasks. [^1] To avoid confusion, we clarify the concept of “pre-training” in this paper. We use this term in its original and broadest sense: any initial training phase that improves model performance on downstream tasks, regardless of the data source (labeled/unlabeled, in-domain/out-of-domain) or methodology (self-supervised, supervised, etc.). This intentionally differs from narrower interpretations that restrict “pre-training” exclusively to large-scale unsupervised learning on massive datasets, as commonly employed in today’s large models. Finally, given the wide diversity of tasks in wireless communication and sensing, designing specialized models for each task can be labor-intensive and inefficient. A unified and generic framework is therefore highly desirable [foundation_survey], and many works [yang2025wirelessgpt, liu2025wifo, jiang2025mimo] have begun exploring the development of foundational models for wireless systems.
@@
To address the aforementioned shortcomings, we propose a method inspired by bidirectional encoder representations from transformers (BERT) [BERT], named CSI-BERT2, which is an improved version of our previous work, CSI-BERT [CSI-BERT]. In CSI-BERT, the masked language model (MLM) approach was first introduced for the pre-training of the CSI recovery model in Wi-Fi sensing, enabling it to learn effectively from limited dataset and improving performance in downstream tasks. Building on this, CSI-BERT2 introduces the mask prediction model (MPM), where we mask the final portion of the CSI sequence and train the model to recover it. By predicting the masked CSI in a single step, our method efficiently avoids the error accumulation problem common in AR models. Additionally, we propose a novel multi-layer perceptron (MLP)-based time embedding method that effectively addresses the temporal information loss problem [transformer-time] inherent in the original Transformer. As illustrated in [transformer-time], the permutation-invariant nature of the attention mechanism leads to temporal information loss, even with positional embeddings. In contrast, our proposed time embedding method explicitly exploits the temporal order, which is critical for CSI sequences, thereby mitigating this issue. This design enables CSI-BERT2 to generalize efficiently in environments with varying sampling rates, and perform robustly even in a zero-shot setting. Finally, considering the varying information content convened across different subcarriers, we introduce an adaptive re-weighting layer (ARL) [arl], which helps to adaptively capture the importance of each subcarrier. Through these architectural enhancements, we successfully address the low performance of the original CSI-BERT in CSI classification tasks. In conclusion, the main contributions of this paper are summarized as follows:
@@ II-A CSI Basics
Following [ma2019wifi], in Wi-Fi communication, CSI characterizes the way Wi-Fi signals propagate from the transmitter (TX) to the receiver (RX). Modern Wi-Fi systems typically employ orthogonal frequency division multiplexing (OFDM) as the modulation technique. OFDM divides the wideband channel into multiple narrowband subcarriers in the frequency domain, each experiencing flat fading for easier processing. In the time domain, data is transmitted as a series of OFDM symbols with a cyclic prefix to prevent interference from multipath delays. The amplitude and phase of CSI are influenced by multipath effects, which include amplitude attenuation and phase shift. Each element of the CSI matrix $H$ is given by:
@@ II-B CSI Prediction
Obtaining accurate CSI matrix in wireless communication remains a challenging task [qi2021acquisition]. The large number of antennas results in a high-dimensional CSI matrix, which in turn makes the channel estimation algorithms difficult to solve. Additionally, when objects in the environment move at high speeds, the channel coherence time shortens considerably, imposing stricter requirements on the channel estimation speed. As illustrated in Fig. 2b, CSI plays a critical role in optimizing wireless communication system performance, making fast and reliable CSI estimation essential. Furthermore, predicting future CSI can provide valuable opportunities for system-level optimizations, such as effective channel compensation.
Recently, some works have explored the use of additional information for CSI prediction. For example, in [9277535], the authors used multimodal sensory data, including received pilots, user position, and historical downlink CSI to predict the current CSI. In [zhang2022predicting], a deep learning-based framework using a three dimensional (3D) CNN architecture effectively predicts downlink CSI by learning temporal, spatial, and frequency correlations from past channel sequences. The authors in [8815557] have proposed an machine learning (ML)-based architecture for estimating channels by using CNN-AR, which achieves gains in prediction quality.
@@ II-C CSI Classification
In wireless sensing, many tasks can be formulated as CSI classification problems, including gesture recognition[widar], fall detection [cai2023falldewideo], and activity recognition [zhang2023ratiofi]. Reviewing CSI basics, human movement in the environment can cause changes in the multipath channel, which in turn changes the extracted CSI matrix. Therefore, by analyzing the pattern of changes in CSI, it is possible to sense human activities.
However, in real-world sensing systems, accurately characterizing human actions are challenging. Complex actions, such as falling, kicking, and squatting, are particularly difficult to model mathematically. Additionally, noise in Wi-Fi extraction devices, including carrier frequency offset and sampling time offset, is hard to completely eliminate. These challenges motivate the use of deep learning techniques for Wi-Fi sensing. For example, in [metafi], the authors have estimated human poses from videos to supervise the training of Wi-Fi CSI. The broad learning system proposed in [fingerprint] aims to achieve fast and accurate Wi-Fi fingerprint localization. Furthermore, [wang20223d] have proposed Wi-Mesh, which utilizes Wi-Fi signals and 2D angle-of-arrival (AoA) estimation to create 3D human meshes using deep learning, showing robustness across various environments and conditions.
@@
To tackle this challenge, our previous work have introduced CSI-BERT [CSI-BERT], a self-supervised learning method designed to capture the internal structure of CSI data and recover lost packets. The recovered CSI was shown to substantially enhance the performance of downstream sensing tasks. In this work, we further optimize the structure of CSI-BERT to improve its ability to capture temporal information, enabling it to successfully perform CSI classification tasks using the discontinuous CSI for training.
@@ III-A Overview
In this section, we will introduce the data preprocessing methods, neural network architecture, pre-training process, and the fine-tuning and inference methods of CSI-BERT2 in turn. As a continuation of CSI-BERT, we first provide a brief introduction to it. (This section will primarily focus on the modified and enhanced components, while more details about CSI-BERT can be found in our previous paper [CSI-BERT].)
Prior to CSI-BERT, several works had already applied BERT [BERT] to wireless sensing tasks such as people localization [BERT-loc1, BERT-loc2] and radio map construction [BERT-map], based on the intuition that the bidirectional encoder of the Transformer [Transformer] can effectively capture the relationships among packets by the self-attention mechanism. However, these studies typically fed discrete signals into the original BERT embedding layer, which leads to significant information loss. In contrast, CSI-BERT first proposed a novel embedding layer tailored to the structure of CSI data, where a multi-layer perceptron (MLP) is used to encode continuous CSI data directly. Additionally, a positional-embedding style time-embedding layer was introduced to capture the temporal information among packets. Inspired by the similarity between packet recovery and masked language modeling (MLM), CSI-BERT successfully achieves high-quality packet recovery through unsupervised learning without labeled data. The recovered CSI data enhances the performance of other sensing models in classification tasks, outperforming traditional interpolation methods that fail to consider the internal relationships among subcarriers.
However, the original CSI-BERT does not fully leverage the advantages of MLM, resulting in subpar performance in downstream tasks. For example, the classification accuracy of CSI-BERT after fine-tuning does not surpass that of the vanilla ResNet18 [Resnet]. This limitation can be attributed to the inadequacies of the embedding layer. To address this issue, CSI-BERT2 introduces an ARL-based spatial encoding module to capture the relationships and importance among subcarriers. Additionally, a MLP-based temporal encoding module is proposed to capture both the order and distance relationships among packets, effectively solving the temporal information loss problem inherent in conventional positional embedding modules, which capture distance relationships without considering order information. Moreover, to fully utilize the learned knowledge during pre-training, we propose a novel fine-tuning method named mask prediction model (MPM) for the CSI prediction task, wherein the last portion of the CSI sequence is randomly masked, and the model is trained to predict these masked values. Compared to conventional AR-based methods, our approach can complete the prediction in one step, thereby eliminating the error accumulation problem.
@@
TABLE I: Comparison between CSI-BERT [CSI-BERT] and CSI-BERT2: Here PE represents the original positional embedding in Transformer, and SFT represents supervised fine-tuning. We utilize underline to highlight the differences between CSI-BERT2 and CSI-BERT.
CSI-BERT [CSI-BERT] CSI-BERT2 Model Structure Spatial Embedding MLP MLP + ARL Temporal Embedding PE PE + MLP Positional Embedding PE PE Backbone BERT BERT Discriminator DANN-based GAN-based Training Method Pre-training MLM MLM Classification Fine-tuning SFT SFT + Random Mask Prediction Fine-tuning $\times$ MPM Application Scenarios CSI Recovery $\checkmark$ $\checkmark$ CSI Classification Single Sampling Rate Various Sampling Rates CSI Prediction $\times$ $\checkmark$